Please answer 2 and 4, and be as detailed as possible. I appreciate you!!

A concrete mixer has been purchased to improve production in IE 2324 class. There are two operators required to operate the concrete mixer at $20/hour, work 8 hours/day and 20 days/month. Utilities required to operate the mixer cost $2,000/month. The production cost of I m concrete is $100 per concrete produced. The revenue to be generated from production of I m of concrete is $500. How many m of concrete need to be sold per month so that the revenue generated can breakeven with the expenditure? Also, if an expected $10,000 profit is expected monthly, how many ‘m of concrete need to be sold? 3. IE 2324 Restaurant, which sells only pepperoni pizza, the expenses per pizza are Variable Costs General Labor $1,390 FlourS0.50 $2,000 Yeast $O.0s Insurance $200 Water $0.01 Advertising $100 Cheese $3.00 Utilities$450 Pepperoni $2.00 Fixed Costs evenve: 1O Rent 4140 How many pizzas does IE 2324 REstaurant need to sell at $10 each to breakeven? If they sold 300 pizzas the first month, did they make a profit or loss? IE2324 class has created a startup and the app is starting to generate revenues! The startup charges a S 29 fee from a user, to have a license. The fixed cost is $ 1,500 (per month) for technical support, and variable cost of S 4 per license sold. It is forecasted that the company will sell 500 licenses per month, for the next year. How many months will you reach the breakeven of your investment? How many licenses do they need to sell to breakeven? 4. a) b)
Answer #1

2. Here we need to find Fixed cost and Variable cost separately

The formula = Fixed cost/(Price-Variable cost)

Fixed cost = Operators cost + Utilities

= (20*8*2*20)+(2000) = 6400+2000 = 8400

Variable cost = 100

Price = 500

Break even = 8400/(500-100) = 21

So, 21 m3 need to be sold to achieve break even

Now, to achieve expected profit of $10000, 46 m3 need to be sold

Q

Fixed cost

VC

TC

TR

Profit

46

8400

100

13000

23000

10000

4.

a) To calculate number of months required break even investment is,

As fixed cost is the investment needs to be covered, in the first month itself the break even is achieved

Month

Q

FC

VC

TC

P

TR

Profit

Cumulative profit

1

500

1500

4

3500

29

14500

11000

11000

So by dividing 500/30 we get roughly 16 licenses per day as sales so,

Day

Q

FC

VC

TC

P

TR

Profit

Cumulative profit

1

16

1500

4

1564

29

464

-1100

-1100

2

16

4

64

29

464

400

-700

3

16

4

64

29

464

400

-300

4

16

4

64

29

464

400

100

5

16

4

64

29

464

400

500

So at 4th day the break even is achieved as profit becomes positive

b) Number of licenses to be sold = Fixed/(Price-Variable cost) = 1500/(29-4) = 60 licenses